OpenStudy (anonymous):

How do you divide: 10 radical 50 over 4 radical 2?

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf \frac{ 10 \sqrt{50} }{4\sqrt{2}} }\) you can brake 50 into 25 and 2. let me know if you need more help with this problem.

OpenStudy (anonymous):

I broke it down to dw:1396836915847:dw|

OpenStudy (solomonzelman):

▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼▼ \(\Huge\color{blue}{ \bf \frac{10\sqrt{50} }{4\sqrt{2}} }\) \(\Huge\color{red}{ \bf \text{_____________________} }\) \(\Huge\color{blue}{ \bf \frac{10\sqrt{25 \times 2} }{4\sqrt{2}} }\) \(\Huge\color{red}{ \bf \text{_____________________} }\) \(\Huge\color{blue}{ \bf \frac{10\sqrt{25} \times \sqrt{2} }{4\sqrt{2}} }\) \(\Huge\color{red}{ \bf \text{_____________________} }\) \(\Huge\color{blue}{ \bf \frac{10\sqrt{5^2} \times \sqrt{2} }{4\sqrt{2}} }\) \(\Huge\color{red}{ \bf \text{_____________________} }\) \(\Huge\color{blue}{ \bf \frac{10\times5 \times \sqrt{2} }{4\sqrt{2}} }\) \(\Huge\color{red}{ \bf \text{_____________________} }\) \(\Huge\color{blue}{ \bf \frac{50\sqrt{2} }{4\sqrt{2}} }\) \(\Huge\color{red}{ \bf \text{_____________________} }\) \(\Huge\color{blue}{ \bf \frac{50 }{4} }\) \(\Huge\color{red}{ \bf \text{_____________________} }\) \(\Huge\color{blue}{ \bf \frac{25 }{2} ~~or,~~~~~12~\frac{1 }{2}}\) \(\Huge\color{red}{ \bf ■■■■■■■■■■■■■■ }\)